# Prompt Attacks Reveal Superficial Knowledge Removal in Unlearning Methods

### v1

@@ Appendix B Evaluation of the Rephrased WMDP-bio Datasets
Table 3: Full experiment results on every rephrasing prompt we tested (logit-based results).
Model Task Accuracy
Zephyr_RMU tinyMMLU 0.6082
Zephyr_RMU wmdp_bio 0.3071
Zephyr_RMU wmdp_cyber 0.2718
Zephyr_RMU wmdp_chem 0.4485
Zephyr_RMU rephrased_english_filler 0.3142
Zephyr_RMU rephrased_hindi_filler 0.3009
Zephyr_RMU rephrased_latin_filler 0.3417
Zephyr_RMU rephrased_conversation 0.3040
Zephyr_RMU rephrased_poem 0.3215
Zephyr_RMU rephrased_replace_with_variables 0.2820
Zephyr_RMU rephrased_technical_terms_removed_1 0.3071
Zephyr_RMU wmdp_bio_rephrased_translated_arabic 0.3087
Zephyr_RMU wmdp_bio_rephrased_translated_bengali 0.2624
Zephyr_RMU wmdp_bio_rephrased_translated_czech 0.2907
Zephyr_RMU rephrased_translated_farsi 0.3016
Zephyr_RMU wmdp_bio_rephrased_translated_french 0.3032
Zephyr_RMU rephrased_translated_german 0.2922
Zephyr_RMU wmdp_bio_rephrased_translated_hindi 0.2899
Zephyr_RMU rephrased_translated_korean 0.2946
Zephyr_RMU wmdp_bio_rephrased_translated_telugu 0.2702
Zephyr_RMU wmdp_bio_rephrased_translated_turkish 0.2859
Zephyr_RMU wmdp_bio_rephrased_translated_vietnamese 0.2720
Llama3-8B-RMU tinyMMLU 0.5595
Llama3-8B-RMU wmdp_bio 0.2773
Llama3-8B-RMU rephrased_english_filler 0.2781
Llama3-8B-RMU rephrased_hindi_filler 0.2797
Llama3-8B-RMU rephrased_latin_filler 0.2757
Llama3-8B-RMU rephrased_conversation 0.2529
Llama3-8B-RMU rephrased_poem 0.2569
Llama3-8B-RMU rephrased_replace_with_variables 0.2781
Llama3-8B-RMU rephrased_technical_terms_removed_1 0.2828
Llama3-8B-RMU rephrased_translated_farsi 0.2789
Llama3-8B-RMU rephrased_translated_german 0.2773
Llama3-8B-RMU rephrased_translated_korean 0.2773
Llama3-TAR-bio tinyMMLU 0.4738
Llama3-TAR-bio wmdp_bio 0.2781
Llama3-TAR-bio rephrased_english_filler 0.3103
Llama3-TAR-bio rephrased_hindi_filler 0.3032
Llama3-TAR-bio rephrased_latin_filler 0.3032
Llama3-TAR-bio rephrased_conversation 0.3032
Llama3-TAR-bio rephrased_poem 0.2868
Llama3-TAR-bio rephrased_replace_with_variables 0.2828
Llama3-TAR-bio rephrased_technical_terms_removed_1 0.2765
Llama3-TAR-bio rephrased_translated_farsi 0.2757
Llama3-TAR-bio rephrased_translated_german 0.2930
Llama3-TAR-bio rephrased_translated_korean 0.2922
Zephyr-7B-ELM tinyMMLU 0.6185
Zephyr-7B-ELM wmdp_bio 0.3016
Zephyr-7B-ELM wmdp_bio_rephrased_english_filler 0.3519
Zephyr-7B-ELM wmdp_bio_rephrased_hindi_filler 0.5507
Zephyr-7B-ELM wmdp_bio_rephrased_latin_filler 0.3778
Zephyr-7B-ELM wmdp_bio_rephrased_conversation 0.2977
Zephyr-7B-ELM wmdp_bio_rephrased_poem 0.2868
Zephyr-7B-ELM wmdp_bio_rephrased_replace_with_variables 0.3252
Zephyr-7B-ELM wmdp_bio_rephrased_technical_terms_removed_1 0.3111
Zephyr-7B-ELM wmdp_bio_rephrased_translated_farsi 0.3621
Zephyr-7B-ELM wmdp_bio_rephrased_translated_german 0.3040
Zephyr-7B-ELM wmdp_bio_rephrased_translated_korean 0.3252
Mistral-7B-ELM tinyMMLU 0.5597
Mistral-7B-ELM wmdp_bio 0.2891
Mistral-7B-ELM wmdp_bio_rephrased_english_filler 0.3064
Mistral-7B-ELM wmdp_bio_rephrased_hindi_filler 0.4721
Mistral-7B-ELM wmdp_bio_rephrased_latin_filler 0.3032
Mistral-7B-ELM wmdp_bio_rephrased_conversation 0.3001
Mistral-7B-ELM wmdp_bio_rephrased_poem 0.3255
Mistral-7B-ELM wmdp_bio_rephrased_replace_with_variables 0.3056
Mistral-7B-ELM wmdp_bio_rephrased_technical_terms_removed_1 0.2875
Mistral-7B-ELM wmdp_bio_rephrased_translated_farsi 0.2844
Mistral-7B-ELM wmdp_bio_rephrased_translated_german 0.2875
Mistral-7B-ELM wmdp_bio_rephrased_translated_korean 0.2954
Llama3-8B-Instruct-ELM tinyMMLU 0.5741
Llama3-8B-Instruct-ELM wmdp_bio 0.3299
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_english_filler 0.3959
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_hindi_filler 0.5373
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_latin_filler 0.3582
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_conversation 0.3472
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_poem 0.3278
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_replace_with_variables 0.3378
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_technical_terms_removed_1 0.2985
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_translated_farsi 0.3040
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_translated_german 0.3221
Llama3-8B-Instruct-ELM wmdp_bio_rephrased_translated_korean 0.3472
Llama3-8B-ELM tinyMMLU 0.6004
Llama3-8B-ELM wmdp_bio 0.3449
Llama3-8B-ELM wmdp_bio_rephrased_english_filler 0.4077
Llama3-8B-ELM wmdp_bio_rephrased_hindi_filler 0.5923
Llama3-8B-ELM wmdp_bio_rephrased_latin_filler 0.3425
Llama3-8B-ELM wmdp_bio_rephrased_conversation 0.4438
Llama3-8B-ELM wmdp_bio_rephrased_poem 0.3381
Llama3-8B-ELM wmdp_bio_rephrased_replace_with_variables 0.2938
Llama3-8B-ELM wmdp_bio_rephrased_technical_terms_removed_1 0.2993
Llama3-8B-ELM wmdp_bio_rephrased_translated_farsi 0.2946
Llama3-8B-ELM wmdp_bio_rephrased_translated_german 0.3024
Llama3-8B-ELM wmdp_bio_rephrased_translated_korean 0.2899
Llama3-8b-instruct-pbj-checkpoint-8 tinyMMLU 0.6118
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio 0.3229
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_conversation 0.3252
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_english_filler 0.3244
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.3221
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.3252
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_poem 0.3522
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.3229
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.3307
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.3009
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_translated_german 0.3056
Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.3032
Llama3-8b-instruct-rr-checkpoint-8 tinyMMLU 0.5852
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio 0.3244
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_conversation 0.2969
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2969
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.3229
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.3268
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_poem 0.3239
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.3181
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.3103
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.3221
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_translated_german 0.3024
Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.3087
Llama3-8b-instruct-tar-checkpoint-8 tinyMMLU 0.4962
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio 0.2710
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_conversation 0.2718
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2954
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.3095
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.2891
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_poem 0.2790
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.2632
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.2561
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.2883
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_translated_german 0.2804
Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.2812
Llama3-8b-instruct-graddiff-checkpoint-8 tinyMMLU 0.4556
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio 0.2742
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_conversation 0.2467
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2498
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.2482
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.2624
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_poem 0.2467
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.2883
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.2812
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.2608
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_translated_german 0.2482
Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.2561
Llama3-8b-instruct-elm-checkpoint-8 tinyMMLU 0.5814
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio 0.3252
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_conversation 0.3519
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_english_filler 0.3873
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.5467
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.3504
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_poem 0.3294
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.3401
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.3150
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.3307
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_translated_german 0.3024
Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.3150
Llama3-8b-instruct-repnoise-checkpoint-8 tinyMMLU 0.3721
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio 0.2529
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_conversation 0.2451
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2459
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.2474
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.2679
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_poem 0.2467
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.2506
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.2522
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.2474
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_translated_german 0.2435
Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.2451
Llama3-8b-instruct-rmu-checkpoint-8 tinyMMLU 0.5329
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio 0.2734
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_conversation 0.2506
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2655
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.2757
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.2836
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_poem 0.2861
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.2875
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.2985
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.2914
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_translated_german 0.2828
Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.2663
Llama3-8b-instruct-rmu-lat-checkpoint-8 tinyMMLU 0.5010
Llama3-8b-instruct-rmu-lat-checkpoint-8 wmdp_bio 0.3001
Llama3-8b-instruct-rmu-lat-checkpoint-8 wmdp_bio_rephrased_conversation 0.2467
Llama3-8b-instruct-rmu-lat-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2828
Llama3-8b-instruct-rmu-lat-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.2765
Llama3-8b-instruct-rmu-lat-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.3111
Llama3-8b-instruct-rmu-lat-checkpoint-8 wmdp_bio_rephrased_poem 0.2782
Llama3-8b-instruct-rmu-lat-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.3174
Llama3-8b-instruct-rmu-lat-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.2969
Llama3-8b-instruct-rmu-lat-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.3071
[…]

### v2

@@ Appendix B Evaluation of the Rephrased WMDP-bio Datasets
Table 3: Full experiment results on every rephrasing prompt we tested on Zephyr_RMU model (logit-based results).
Model Task Accuracy Llama3-8B-RMU tinyMMLU 0.5595 Llama3-8B-RMU wmdp_bio 0.2773 Llama3-8B-RMU rephrased_english_filler 0.2781 Llama3-8B-RMU rephrased_hindi_filler 0.2797 Llama3-8B-RMU rephrased_latin_filler 0.2757 Llama3-8B-RMU rephrased_conversation 0.2529 Llama3-8B-RMU rephrased_poem 0.2569 Llama3-8B-RMU rephrased_replace_with_variables 0.2781 Llama3-8B-RMU rephrased_technical_terms_removed_1 0.2828 Llama3-8B-RMU rephrased_translated_farsi 0.2789 Llama3-8B-RMU rephrased_translated_german 0.2773 Llama3-8B-RMU rephrased_translated_korean 0.2773
Table 4: Full experiment results on every rephrasing prompt we tested on Llama3-8B-RMU model (logit-based results).
Model Task Accuracy Llama3-TAR-bio tinyMMLU 0.4738 Llama3-TAR-bio wmdp_bio 0.2781 Llama3-TAR-bio rephrased_english_filler 0.3103 Llama3-TAR-bio rephrased_hindi_filler 0.3032 Llama3-TAR-bio rephrased_latin_filler 0.3032 Llama3-TAR-bio rephrased_conversation 0.3032 Llama3-TAR-bio rephrased_poem 0.2868 Llama3-TAR-bio rephrased_replace_with_variables 0.2828 Llama3-TAR-bio rephrased_technical_terms_removed_1 0.2765 Llama3-TAR-bio rephrased_translated_farsi 0.2757 Llama3-TAR-bio rephrased_translated_german 0.2930 Llama3-TAR-bio rephrased_translated_korean 0.2922
Table 5: Full experiment results on every rephrasing prompt we tested on Llama3-TAR-bio model (logit-based results).
Model Task Accuracy Zephyr-7B-ELM tinyMMLU 0.6185 Zephyr-7B-ELM wmdp_bio 0.3016 Zephyr-7B-ELM wmdp_bio_rephrased_english_filler 0.3519 Zephyr-7B-ELM wmdp_bio_rephrased_hindi_filler 0.5507 Zephyr-7B-ELM wmdp_bio_rephrased_latin_filler 0.3778 Zephyr-7B-ELM wmdp_bio_rephrased_conversation 0.2977 Zephyr-7B-ELM wmdp_bio_rephrased_poem 0.2868 Zephyr-7B-ELM wmdp_bio_rephrased_replace_with_variables 0.3252 Zephyr-7B-ELM wmdp_bio_rephrased_technical_terms_removed_1 0.3111 Zephyr-7B-ELM wmdp_bio_rephrased_translated_farsi 0.3621 Zephyr-7B-ELM wmdp_bio_rephrased_translated_german 0.3040 Zephyr-7B-ELM wmdp_bio_rephrased_translated_korean 0.3252
Table 6: Full experiment results on every rephrasing prompt we tested on Zephyr-7B-ELM model (logit-based results).
Model Task Accuracy Mistral-7B-ELM tinyMMLU 0.5597 Mistral-7B-ELM wmdp_bio 0.2891 Mistral-7B-ELM wmdp_bio_rephrased_english_filler 0.3064 Mistral-7B-ELM wmdp_bio_rephrased_hindi_filler 0.4721 Mistral-7B-ELM wmdp_bio_rephrased_latin_filler 0.3032 Mistral-7B-ELM wmdp_bio_rephrased_conversation 0.3001 Mistral-7B-ELM wmdp_bio_rephrased_poem 0.3255 Mistral-7B-ELM wmdp_bio_rephrased_replace_with_variables 0.3056 Mistral-7B-ELM wmdp_bio_rephrased_technical_terms_removed_1 0.2875 Mistral-7B-ELM wmdp_bio_rephrased_translated_farsi 0.2844 Mistral-7B-ELM wmdp_bio_rephrased_translated_german 0.2875 Mistral-7B-ELM wmdp_bio_rephrased_translated_korean 0.2954
Table 7: Full experiment results on every rephrasing prompt we tested on Mistral-7B-ELM model (logit-based results).
Model Task Accuracy Llama3-8B-Instruct-ELM tinyMMLU 0.5741 Llama3-8B-Instruct-ELM wmdp_bio 0.3299 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_english_filler 0.3959 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_hindi_filler 0.5373 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_latin_filler 0.3582 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_conversation 0.3472 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_poem 0.3278 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_replace_with_variables 0.3378 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_technical_terms_removed_1 0.2985 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_translated_farsi 0.3040 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_translated_german 0.3221 Llama3-8B-Instruct-ELM wmdp_bio_rephrased_translated_korean 0.3472
Table 8: Full experiment results on every rephrasing prompt we tested on Llama-8B-Instruct-ELM model (logit-based results).
Model Task Accuracy Llama3-8B-ELM tinyMMLU 0.6004 Llama3-8B-ELM wmdp_bio 0.3449 Llama3-8B-ELM wmdp_bio_rephrased_english_filler 0.4077 Llama3-8B-ELM wmdp_bio_rephrased_hindi_filler 0.5923 Llama3-8B-ELM wmdp_bio_rephrased_latin_filler 0.3425 Llama3-8B-ELM wmdp_bio_rephrased_conversation 0.4438 Llama3-8B-ELM wmdp_bio_rephrased_poem 0.3381 Llama3-8B-ELM wmdp_bio_rephrased_replace_with_variables 0.2938 Llama3-8B-ELM wmdp_bio_rephrased_technical_terms_removed_1 0.2993 Llama3-8B-ELM wmdp_bio_rephrased_translated_farsi 0.2946 Llama3-8B-ELM wmdp_bio_rephrased_translated_german 0.3024 Llama3-8B-ELM wmdp_bio_rephrased_translated_korean 0.2899
Table 9: Full experiment results on every rephrasing prompt we tested on Llama-8B-ELM model (logit-based results).
Model Task Accuracy Llama3-8b-instruct-pbj-checkpoint-8 tinyMMLU 0.6118 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio 0.3229 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_conversation 0.3252 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_english_filler 0.3244 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.3221 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.3252 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_poem 0.3522 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.3229 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.3307 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.3009 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_translated_german 0.3056 Llama3-8b-instruct-pbj-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.3032
Table 10: Full experiment results on every rephrasing prompt we tested on Llama-8B-Instruct-PBJ model (logit-based results).
Model Task Accuracy Llama3-8b-instruct-rr-checkpoint-8 tinyMMLU 0.5852 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio 0.3244 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_conversation 0.2969 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2969 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.3229 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.3268 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_poem 0.3239 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.3181 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.3103 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.3221 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_translated_german 0.3024 Llama3-8b-instruct-rr-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.3087
Table 11: Full experiment results on every rephrasing prompt we tested on Llama-8B-Instruct-$R$ model (logit-based results).
Model Task Accuracy Llama3-8b-instruct-tar-checkpoint-8 tinyMMLU 0.4962 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio 0.2710 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_conversation 0.2718 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2954 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.3095 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.2891 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_poem 0.2790 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.2632 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.2561 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.2883 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_translated_german 0.2804 Llama3-8b-instruct-tar-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.2812
Table 12: Full experiment results on every rephrasing prompt we tested on Llama-8B-Instruct-TAR model (logit-based results).
Model Task Accuracy Llama3-8b-instruct-graddiff-checkpoint-8 tinyMMLU 0.4556 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio 0.2742 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_conversation 0.2467 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2498 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.2482 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.2624 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_poem 0.2467 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.2883 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.2812 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.2608 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_translated_german 0.2482 Llama3-8b-instruct-graddiff-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.2561
Table 13: Full experiment results on every rephrasing prompt we tested on Llama-8B-Instruct-GradDiff model (logit-based results).
Model Task Accuracy Llama3-8b-instruct-elm-checkpoint-8 tinyMMLU 0.5814 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio 0.3252 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_conversation 0.3519 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_english_filler 0.3873 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.5467 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.3504 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_poem 0.3294 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.3401 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.3150 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.3307 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_translated_german 0.3024 Llama3-8b-instruct-elm-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.3150
Table 14: Full experiment results on every rephrasing prompt we tested on Llama-8B-Instruct-ELM model (logit-based results).
Model Task Accuracy Llama3-8b-instruct-repnoise-checkpoint-8 tinyMMLU 0.3721 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio 0.2529 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_conversation 0.2451 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2459 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.2474 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.2679 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_poem 0.2467 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.2506 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.2522 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.2474 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_translated_german 0.2435 Llama3-8b-instruct-repnoise-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.2451
Table 15: Full experiment results on every rephrasing prompt we tested on Llama-8B-Instruct-RepNoise model (logit-based results).
Model Task Accuracy Llama3-8b-instruct-rmu-checkpoint-8 tinyMMLU 0.5329 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio 0.2734 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_conversation 0.2506 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_english_filler 0.2655 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_hindi_filler 0.2757 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_latin_filler 0.2836 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_poem 0.2861 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_replace_with_variables 0.2875 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_technical_terms_removed_1 0.2985 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_translated_farsi 0.2914 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_translated_german 0.2828 Llama3-8b-instruct-rmu-checkpoint-8 wmdp_bio_rephrased_translated_korean 0.2663
Table 16: Full experiment results on every rephrasing prompt we tested on Llama-8B-Instruct-RMU model (logit-based results).
[…]